A purification system with recycle is used to recover the solvent DMF from a 100 lb/hr waste gas stream containing 55% DMF in air. The product is to have only 10% DMF. The purification unit was designed to remove two thirds of the DMF in the combined feed to the unit. (Compositions in mass fraction) Determine: 1. Calculate the recycle fraction. 2. Calculate the composition and flow rates of all the streams.
Answer #1

Let R =recycle, Since the products is gas and gases are miscible in all proportions, the recycle also will have same composition as product

Fresh feed = 100 lb/hr contains 45% air. Mass of air = 45 lb/hr

Product contains 10% DMF.   Let P =product

P*0.9 =45 and hence P= 45/0.9= 61 lb/hr

Hecnce DMF removed = 100-61= 39 lb/hr

Feed to treatment plant = (100+R)

DMF to treatment plant = 100*0.55+R*0.1

DMF removed = 2/3

Hecne   DMF removed = 2/3(100*0.55+R*0.1)= 39

55+R*0.1= 39*3/2=58.5

R= 3.5/0.1= 35 lb/hr

Fraction recycled= 35/(35+61)=0.3645

Flow rate of feed to treatment plant = 100+35= 135

Composition : DMF : (100*0.55+35*0.1)/135 = 0.433 and air =1-0.433= 0.567